simplify:
(x^3+y^3)/(x+y)
thanks:)

Answer by Nate(3500)

(x)^3 + (y)^3 = (x + y)(x^2 - yx + y^2)
so:
(x + y)(x^2 - yx + y^2)/(x + y) = x^2 - yx + y^2